OpenStudy (anonymous):

Same question but just CHECK IT Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4). Do not use spaces and enter fractions as improper reduced fractions when needed. y+4=-3x+9 this is my answer but its saying its wrong Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4). Do not use spaces and enter fractions as improper reduced fractions when needed. y=0x+4 and this one too also

OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of -3/4 and passes through the point (4, 5). Do not use spaces and enter fractions as improper reduced fractions when needed. y-5=-3/4x-4

OpenStudy (ash2326):

You need to simplify i \[y+4=-3x+9\] Write it as \[y=-3x+9-4=>y=-3x+5\]

OpenStudy (anonymous):

confused question

OpenStudy (ash2326):

Second one is just \[y=0x+4=>y=4\]

OpenStudy (ash2326):

For third one also bring the -5 from right side to left side \[y=\frac{-3}{4}x+5-4\] We'll get \[y=\frac{-3}{4}x+1\]

OpenStudy (anonymous):

thank you Ash so much... @ash2326

OpenStudy (ash2326):

You got them right?

OpenStudy (anonymous):

nope not yet.. i will submit it very soon cuz i really need to be done

OpenStudy (ash2326):

cool:)